OpenStudy (anonymous):

find the distance between the points (2,-3) (5,-4)

OpenStudy (anonymous):

Use this formula \[distance = \sqrt{(x_2-x_1)^2 + (y_2-y_1)^2}\] where \[(x_1,y_1) \space and \space (x_2,y_2) are \space the \space two \space points \space given\] Now use this to find your answer

OpenStudy (anonymous):

so the distance is 10 ?

OpenStudy (anonymous):

Not exactly. It should be \[\sqrt{10}\]

OpenStudy (anonymous):

so does that mean i sqaure root it or it just stays that way?

OpenStudy (anonymous):

See the formula again. The square root is present. You have not accounted the square root in the formula while working out the sum

OpenStudy (anonymous):

so do i take the swuare root of 10? cuz i just get 10.

OpenStudy (anonymous):

\[\LARGE d=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\] \[\LARGE d=\sqrt{(5-2)^2+(-4-3 )^2}\] \[\LARGE d=\sqrt{(3)^2+(-7 )^2}\] \[\LARGE d=\sqrt{9+49}\] \[\LARGE d=\sqrt{58}\]

OpenStudy (anonymous):

@Kreshnik It should be - 4 +3 not -4-3. Check again

OpenStudy (anonymous):

@shivam_bhalla You're right.. thanks for correcting me..

OpenStudy (anonymous):

\[\LARGE d=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2 }\] \[\LARGE d=\sqrt{(5-2)^2+(-4-(-3))^2 }\] \[\LARGE d=\sqrt{(3)^2+(-1)^2 }\] \[\LARGE d=\sqrt{9+1 }\] \[\LARGE d=\sqrt{10 }\] :)

OpenStudy (anonymous):

@RiyaMoon , I hope you now understand how we get sqrt(10) as the answer .

OpenStudy (anonymous):

what the heck im confused......i just said square root 10 was the anwser and you told me it wasnt...

OpenStudy (anonymous):

@RiyaMoon , you told me "so the distance is 10 ?"

OpenStudy (anonymous):

yea then after you told me to look it over and i said so i have to take the square root of 10 which is 10.

OpenStudy (anonymous):

@RiyaMoon , you told me "so do i take the swuare root of 10? cuz i just get 10." ??

OpenStudy (anonymous):

nvm....

OpenStudy (anonymous):

Ok. :D

OpenStudy (anonymous):

hey guys, don't argue... the best way to avoid confusions is explaining it ;)

OpenStudy (anonymous):

@Kreshnik , :D :D .